OpenStudy (anonymous):

Evaluate (solve)... (8/27)to the power of 2/3 The answer is 4/9 ... how? Explanations, please?

OpenStudy (anonymous):

type into ur calculator 8 / 27....

OpenStudy (anonymous):

I have to solve this without a calculator :(

OpenStudy (anonymous):

oh crap tht suck m sry...idk how to do tht..@1210 :(

OpenStudy (anonymous):

Its ok, thanks though :)

OpenStudy (anonymous):

it is easy to solve (8/27)^2/3 = cube root of(8^2/27^2)=cube root of( 8 * 8/ 27 * 27)=2 * 2/3*3=4/9 Ans